OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) 1 –7 2 –10 3 –13 4 –16 A. f(x) = -1/3x – 4 B. f(x) = 3x + 4 C. f(x) = –3x – 4 D. f(x) = x – 7

OpenStudy (anonymous):

since it is linear, you can use two points to find the equation of the line in y=mx+b format. m, the slope = (y2-y1)/(x2-x1) = (-16--7)/(4-1)=-9/3=-3. b is the y intercept. you can find b by placing one of the x,y points into the equation and solve for b y=-3x+b, b=y+3x = -7+3*1=-4, so y=-3x-4

OpenStudy (anonymous):

so is c

OpenStudy (loser66):

yeap, baby